let f(x)=ax^3 + bx^2 + cx + dsinx.Find d condition dat f(x) is always one-one function.

Grade:

1 Answers

Aman Bansal
12 years ago

Dear MAnuj,

 f(x)= ax3+ bx2+ cx+dsinx

d f (x)/dx=3ax^2 + 2bx + c + dcosx <0 (say)

i.e. 3ax^2 + 2bx + c + d<0   (as max. of cosx =1)

and b^2 -4ac<0 (for im roots)

u get : b^2 < 3a(c+d)

but if u hd taken 3ax^2 + 2bx + c + dcosx >0

then take 3ax^2 + 2bx + c - d>0 

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert